# Generalizing Verifiable Instruction Following

### v2

@@ Abstract
A crucial factor for successful human and AI interaction is the ability of language models or chatbots to follow human instructions precisely. A common feature of instructions are output constraints like “only answer with yes or no" or “mention the word ‘abrakadabra’ at least 3 times" that the user adds to craft a more useful answer. Even today’s strongest models struggle with fulfilling such constraints. We find that most models strongly overfit on a small set of verifiable constraints from the benchmarks that test these abilities, a skill called precise instruction following, and are not able to generalize well to unseen output constraints. We introduce a new benchmark, IFBench, to evaluate precise instruction following generalization on 58 new, diverse, and challenging verifiable out-of-domain constraints. In addition, we perform an extensive analysis of how and on what data models can be trained to improve precise instruction following generalization. Specifically, we carefully design constraint verification modules and show that reinforcement learning with verifiable rewards (RLVR) significantly improves instruction following. In addition to IFBench, we release 29 additional new hand-annotated training constraints and verification functions, RLVR training prompts, and code.
@@ 1 Introduction
Following instructions exactly is a crucial skill for a language model to have, in order for it to generate a useful output that corresponds to the entirety of a user’s specifications and not just the general topic. In particular, instructions often include output constraints that specify length, format and content. A model’s ability to follow constraints in instructions is evaluated on precise instruction following benchmarks with verifiable constraints, with IFEval zhou2023instruction being the most popular benchmark (and it has quickly saturated, with many leading models scoring 80+at as small as 2B parameters dubey2024llama ; lambert2024t ; team2025gemma . [^1] Technically not every constraint passed by users is simple to verify with software, but these evaluations focus on implementable variants for ease of iteration and improvement. This evaluation benchmark consists of a set of 25 constraint templates, which can all be automatically verified using short python functions. Most models strongly overfit to this small set of constraints.
@@
To facilitate experimenting with the generalization of precise IF, we create 29 training constraints by manually curating useful and challenging constraints and verification functions, some of which representative of real-world chatbot usage, and others inspired by the core capabilities we desire our models to have. We find that increasing the number and variety of training constraints improves IF generalization. In addition to creating distinct training constraints, we explore new methods for inducing strong IF performance by appending combinations of constraints and using wider constraint variable ranges for the training prompts in the RL stage, rewarding models for following more complex instructions. For example, an instruction could be combined with a length constraint, a formatting constraint and a constraint asking to include specific words.
@@
    A new, unseen and challenging precise instruction following benchmark, IFBench  [^2] Code for IFBench is available here: https://github.com/allenai/IFBench., with 58 new constraints and corresponding verification functions. With an investigation into the generalization abilities of LLMs for following constraints, we find that leading language models such as Qwen3-32B or Claude 4 Sonnet score below 50% showcasing the opportunity for improvement in precise IF.
@@
    New methods of RLVR training for precise instruction following (IF-RLVR) by interleaving multiple constraints per prompt or mixing verifiable and preference rewards. With our new training techniques we improve the IFeval scores of a Tülu-3-8B model from 82.4 to 92.2, and the IFBench scores from 28.9 to 45.9. Similarly, IF-RLVR increases a Qwen 2.5 7b base model to scores of 87.8 (IFEval) and 54.7 (IFBench), and we also show that our approach is effective for models from the OLMo family.
@@
Figure 1: Model performance on IFEval and IFBench. Left Models: out-of-the-box performance. Right Models: after IF-RLVR training.
@@ 2 IFBench & IFTrain: Measuring and Training Precise IF
We introduce a new benchmark, IFBench, and paradigm to evaluate the generalizability of methods addressing precise instruction following. We define a taxonomy of constraint templates, which we split into training and test constraints to prevent contamination. The new constraints we introduce were created manually – sourced by collecting feedback from LM users beyond the authors on the types of constraints they have tried with models, or manually written to cover core IF skills. Then, we filtered constraints for the benchmark to those that can be easily paired with a verification function written in Python, making for reproducible evaluation and training tools. The full list of new constraints can be found in Appendix A for evaluation and Appendix B for training constraints.
@@
consists of 58 new verifiable constraints that go beyond the 25 constraints included in IFEval zhou2023instruction . To create the final test prompts, we add instantiated constraints to unseen, i.e. held out from release, prompts from WildChat zhao2024wildchat . By combining unseen prompts with unseen constraints, we prevent accidental train-test contamination and can appropriately evaluate language models’ abilities to generalize on the task of precise instruction following. Every instance went through a human annotation process to verify the prompt and constraint compatibility (i.e. a coding related constraint, for example, does not fit with a prompt asking for a summary). These constraints cover 7 different broader categories: count, ratio, words, sentence, format, custom, copy. These categories cover a broad range of sub-skills, such as a model’s ability to copy parts of the input prompt into the output. The final benchmark consists of 300 prompts. When curating the benchmark, another focus was to include challenging constraints, such as Maintain a 2:1 ratio of declarative to interrogative sentences. Similarly to the original IFEval, we compute both strict and loose accuracy, where the strict accuracy verifies if the constraint is followed correctly, and the loose accuracy additionally cleans the model’s output by removing first/last lines and certain font modifiers. We evaluate constraint following abilities in two different settings:
@@ 3 IF-RLVR
Reinforcement Learning with Verifiable Rewards (RLVR) lambert2024t can be applied to the precise instruction following task, as each constraint can be verified with a function. We use GRPO shao2024deepseekmath to optimize the following objective:
Specifically, we train a policy with GRPO and outcome supervision, where each output is scored according to wether or not the constraint has been correctly fulfilled.
@@
IF-RLVR training works for base models, using a special chat template (see Appendix D), and for post-trained models, using their own chat templates.
Experimental Setup
We experiment with the following base policies: Llama-3.1-Tulu-3-8B-DPO, Llama-3.1-8B, Qwen2.5-7B, Qwen2.5-7B-instruct, OLMo2, OLMo2-instruct. We train using the GRPO implementation in open-instruct Ivison2023CamelsIA ; lambert2024t , with the following hyperparameters: max_token_length$=$2048, temperature$=$1, learning rate$=$5e-7, 16 samples per prompt, 8 GPUs and a local mini-batch size of 32. Training took on average 1 day for 2000 steps. Base models with a reasoning chat template are trained with a max token length of 10240, a beta of 0, and a temperature of 1.
@@ 4.1 Training on Multiple Constraints
Figure 4: Training on IFTrain (ood) + n constraints (in-domain) from IFEval. (Tülu-DPO policy)
@@ 4.3 Changing the range of constraint variables
Most of the constraint templates contain variables which can be instantiated with different values. In your response, all lowercase words should appear at most N times., for example, has the variable $N$ which could in theory be any number. For both the IFEval and the IFBench benchmarks, variables are instantiated for each instruction instance from a fixed range of options. Another type of generalization to assess is whether a model trained on the same constraints, but different variable ranges, can still apply the skill to unseen ranges. To evaluate this, we chose variable ranges for training that are disjoint from the test ranges. For the constraint Your response should contain at most num_sentences sentences., for example, we sample a value between 20 and 40 for train and a value between 1 and 20 for test. Specifically we experiment with three different settings: different range, where every train variable is filled from a range that is disjoint from the test range; wider range, where every train variable is filled from a range that includes and extends the test range; same range, where variables are filled with the same range for train and test. Fig. 5 shows performance on IFEval for different steps of IF-RLVR training. Training on a different variable range consistently scores lower than the other two setups. Interestingly, though, training on a wider variable range, performs comparably and often even better than training on the same range. This suggests that training on a diverse set of constraint variables improves generalization for in-domain constraints performance.
@@ 4.4 Removing Categories
Most verifiable constraints fall under broader constraint type categories. IFEval, for example, has 9 different categories, such as length constraints or detectable format. To investigate how training on a set of categories generalizes to an unseen categories, we experiment with training on a leave-one-out set of categories, iteratively removing one of the following classes: change cases, detectable format, length constraints, keywords.
Removing the constraints from the length constraint and the keywords categories harms IFEval performance the most, while removing constraints from the change cases and detectable format categories barely affect performance with the model achieving an IFEval accuracy of 89.65 (see Fig. 6).
@@ 4.5 Teaching the basic units of precise IF
We designed the new training constraints so that they would cover IF skills models are currently lacking in, such as copying from the input, counting, and formatting. We find that GRPO training on our new constraints shows targeted improvements in all these areas. As seen in Table 2, our final two models (from base vs. from instruct) improve over the base model (in this case a DPO trained model) in all categories, such as in counting, inserting the right amount of keywords, formatting and in following length constraints. Most of the IFEval categories seem saturated with IF-RLVR training (performance >90>90> 90), while the IFBench categories leave room for improvement, such as the words and sentence categories.
Comparing our final models to other post-trained (SFT, DPO, RLVR) models in Table 3, we find that our approach and model results in the best in- and out-of-domain instruction following performance on both IFEval and IFBench. We also see that targeted RLVR training for IF slightly harms other downstream evaluations, such as AlpacaEval 2, while staying comparable on others, such as GSM8K, MMLU and BBH. We therefore suggest, for future work, to investigate how to combine precise IF RLVR with other types of rewards for other tasks such as math or chat. Note that the performance on other benchmarks is low for IF-RLVR Qwen2.5, as the policy is a base model that has not been post-trained.
@@ 4.6 DPO
We generate prompts using the approach described in Section 3, with up to 5 constraints per prompt, and sample completions from 5 different models (Tülu-3-70B, Qwen-72b, Llama-3.1-405b, Llama3-8b, Yi-34B-Chat). For each constraint in a prompt a completion is verified on whether it fulfills the constraint. We construct preference pairs by sampling, for each instruction, a completion that satisfies all constraints (chosen) and a completion that does not satisfy at least one constraint (rejected). Table 4, shows how most models struggle with precisely following a lot of our prompts, and even when relaxing the requirement and counting cases where models got not more than 1 constraint wrong, the percentages are less than half. The model that stands out is Qwen-72B. Out of all the instances that models get completely correct, 54% have only 1 constraint and only 2% have 5 constraints. We also find that most LLMs get the same easy instances right and the same hard instances wrong, which makes the creation of preference pairs more difficult. This further highlights the benefits of RLVR, where we can get ground truth signal for both easy and hard prompts, with an unlimited amount of constraints. We end up with a set of prompts and chosen/rejected pairs, called strict, where chosen completions get all constraints correct (31751 prompts).
@@ 4.7 RLVR for IF from Base
We compare IF-RLVR training using policy models that have gone through post-training already (SFT and DPO), with IF-RLVR training from base models, as it has been shown that it is also possible to perform RLVR training on a base model, leading to good math and reasoning performance shao2024deepseekmath ; wang2025reinforcement ; yang2025qwen3 .
Specifically, we RLVR train three base models for precise IF, llama3.1-8b, Qwen2.5-7B and Qwen3-8B, and their instruct counterparts. To encourage reasoning we use a special chat template (see Appendix D). The models are trained with a max token length of 10240, a beta of 0, and a temperature of 1.
@@ 4.8 RLVR for Multi-turn IF
We experiment with comparing single-turn RLVR runs (IF-RLVR-single) to multi-turn RLVR training (IF-RLVR-multi), and to training on a mix of both types of data (IF-RLVR-mix). We run experiments on Qwen2.7-7b base and instruct.
Table 7: Training on single turn data, multi turn data, and a mix. Evaluated on IFEval (IFE.) constraints and IFBench (IFB.) constraints.
trained on single turn trained on multi turn trained on a mix
single turn multi turn single turn multi turn single turn multi turn
@@
IF-RLVR-multi mostly leads to an improved performance on the multiturn setup of IFBench, compared to IF-RLVR-single, while harming the singleturn performance. IF-RLVR-mix harms singleturn performance less, and sometimes even helps it, while reaching a comparable multiturn performance.
@@ 5 Reward Hacking and the Instruction Hierarchy
To understand the trade-off between challenging constraints and general response quality we contrast IFBench accuracy with a general LLM-as-a-judge prompt-completion rating (zheng2023judging, ). Using GPT4.1 as the judge, we prompt [^3] See Appendix C for details. it to score how well a completion answers a prompt without the constraint. We score the IFEval and IFBench completions from our RLVR trained model and from the base policy before RLVR training. Completions from the base policy are scored higher by the LLM-as-judge than completions from the IF-RLVR trained model, for both IFEval and IFBench prompts with constraints removed (see Figure 7). This indicates that the base policy models are better at following general instructions, while IF-RLVR trained models are better at following the constraints. The verifiable accuracy, though, is on average higher for the model that went through RLVR training. The verifiable rewards teach the model to prioritize the constraints and in the following section we propose an approach to soften this preference.
We also analyze IFBench completions of existing models: claude-3-7-sonnet, Gemini, Qwen2.5-72B-Instruct, and tulu3-70B. We find that tulu3-70B’s IF accuracy correlates the most with the LLM-as-judge scores and that Qwen2.5-72B-Instruct’s IF accuracy is the most negatively correlated with LLM-as-judge scores, out of this set of models. This indicates that Qwen2.5 tends to focus on the constraints rather than the general instruction.
Mitigating Reward Hacking
While GRPO training with verifiable rewards for precise IF is great at teaching LLMs to follow output constraints, it can sometimes result in models that over-prioritize the constraint over the full instruction. An example of such an output is given in Figure 8. This could also be called over-optimization. We propose adding a general reward model (RM) signal to the verifiable reward. The intuition is that while the verifiable reward checks for the adherence to the output constraint, the general reward model provides signal for whether the response answers the prompt. We apply the reward from the reward model to all generations that received a verifiable reward >0>0> 0, as follows: For a batch of data, we assign the final reward, $F_{i}$ to that instance corresponding to conditions of the verifiable reward, $V_{i}$, and the reward model score, $S_{i}$:
[…]

### v3

@@ Abstract
A crucial factor for successful human and AI interaction is the ability of language models or chatbots to follow human instructions precisely. A common feature of instructions are output constraints like “only answer with yes or no" or “mention the word ‘abracadabra’ at least 3 times" that the user adds to craft a more useful answer. Even today’s strongest models struggle with fulfilling such constraints. We find that most models strongly overfit on a small set of verifiable constraints from the benchmarks that test these abilities, a skill called precise instruction following, and are not able to generalize well to unseen output constraints. We introduce a new benchmark, IFBench, to evaluate precise instruction following generalization on 58 new, diverse, and challenging verifiable out-of-domain constraints. In addition, we perform an extensive analysis of how and on what data models can be trained to improve precise instruction following generalization. Specifically, we carefully design constraint verification modules and show that reinforcement learning with verifiable rewards (RLVR) significantly improves instruction following. In addition to IFBench, we release 29 additional new hand-annotated training constraints and verification functions, RLVR training prompts, and code.
@@ 1 Introduction
Following instructions exactly is a crucial skill for a language model to have, in order for it to generate a useful output that corresponds to the entirety of a user’s specifications and not just the general topic. In particular, instructions often include output constraints that specify length, format and content. A model’s ability to follow constraints in instructions is evaluated on precise instruction following benchmarks with verifiable constraints, with IFEval zhou2023instruction being the most popular benchmark (and it has quickly saturated, with many leading models scoring 80+% at as small as 2B parameters dubey2024llama; lambert2024t; team2025gemma. [^1] Technically not every constraint passed by users is simple to verify with software, but these evaluations focus on implementable variants for ease of iteration and improvement. This evaluation benchmark consists of a set of 25 constraint templates, which can all be automatically verified using short python functions. Most models strongly overfit to this small set of constraints.
@@
To facilitate experimenting with the generalization of precise IF, we create 29 training constraints by manually curating useful and challenging constraints and verification functions, some of which are representative of real-world chatbot usage, while others are inspired by the core capabilities we desire our models to have. We find that increasing the number and variety of training constraints improves IF generalization. In addition to creating distinct training constraints, we explore new methods for inducing strong IF performance by appending combinations of constraints and using wider constraint variable ranges for the training prompts in the RL stage, rewarding models for following more complex instructions. For example, an instruction could be combined with a length constraint, a formatting constraint and a constraint asking to include specific words.
@@
    A new, unseen and challenging precise instruction following benchmark, IFBench  [^2] Code for IFBench is available here: https://github.com/allenai/IFBench., with 58 new constraints and corresponding verification functions. With an investigation into the generalization abilities of LLMs for following constraints, we find that leading LMs such as Qwen3-32B or Claude 4 Sonnet score below 50% showcasing the opportunity for improvement in precise IF.
@@
    New methods of RLVR training for precise instruction following (IF-RLVR) by interleaving multiple constraints per prompt or mixing verifiable and preference rewards. With our new training techniques we improve the IFeval scores of a Tülu-3-8B model from 82.4 to 92.2, and the IFBench scores from 28.9 to 45.9. Similarly, IF-RLVR increases a Qwen-2.5-7B base model to scores of 87.8 (IFEval) and 54.7 (IFBench), and we also show that our approach is effective for models from the OLMo family.
@@
Figure 1: Model performance on IFEval and IFBench (single-turn). Left Models: out-of-the-box performance. Right Models: after IF-RLVR training. IFBench has either 1 or 2 constraints per instruction.
@@ 2 IFBench & IFTrain: Measuring and Training Precise IF
We introduce a new benchmark, IFBench, and paradigm to evaluate the generalizability of methods addressing precise instruction following. We define a taxonomy of constraint templates, which we split into training and test constraints to prevent contamination. The new constraints we introduce were created manually – sourced by collecting feedback from LM users beyond the authors on the types of constraints they have tried with models, or manually written to cover core IF skills. Then, we filtered constraints for the benchmark to those that can be easily paired with a verification function written in Python, making for reproducible evaluation and training tools. The full list of new constraints can be found in Appendix A for evaluation and Appendix B for training constraints, with an analysis in Appendix D.
@@
consists of 58 new verifiable constraints that go beyond the 25 constraints included in IFEval zhou2023instruction. To create the final test prompts, we add instantiated constraints to unseen, i.e. held out from release, prompts from WildChat zhao2024wildchat. By combining unseen prompts with unseen constraints, we prevent accidental train-test contamination and can appropriately evaluate language models’ abilities to generalize on the task of precise instruction following. Every instance went through a human annotation process to verify the prompt and constraint compatibility (i.e. a coding related constraint, for example, does not fit with a prompt asking for a summary). These constraints cover 7 different broader categories: count, ratio, words, sentence, format, custom, copy. These categories cover a broad range of sub-skills, such as a model’s ability to copy parts of the input prompt into the output. The final benchmark consists of 300 prompts. When curating the benchmark, another focus was to include challenging constraints, such as Maintain a 2:1 ratio of declarative to interrogative sentences. Similarly to the original IFEval, we compute both strict and loose accuracy, where the strict accuracy verifies if the constraint is followed correctly, and the loose accuracy additionally cleans the model’s output by removing first/last lines and certain font modifiers. Each instruction in the benchmark contains either one or two constraints. We evaluate constraint following abilities in two different settings:
@@ 3 IF-RLVR
Reinforcement Learning with Verifiable Rewards (RLVR) lambert2024t can be applied to the precise instruction following task, as each constraint can be verified with a function. Specifically, we train a policy with GRPO shao2024deepseekmath and outcome supervision, where each output is scored according to wether or not the constraint has been correctly fulfilled.
@@
Where the reward multiplier and the reward weights are hyperparameters, generally set to 1, except if one wants to up or down-weight certain rewards. IF-RLVR training works for base models, using a special chat template (see Appendix F), and for post-trained models, using their own chat templates.
Experimental & Evaluation Setup
We experiment with the following base policies: Llama-3.1-Tulu-3-8B-DPO, Llama-3.1-8B, Qwen2.5-7B, Qwen2.5-7B-instruct, OLMo2, OLMo2-instruct. We train using the GRPO implementation in open-instruct Ivison2023CamelsIA; lambert2024t, with the following hyperparameters: max_token_length==2048, temperature==1, learning rate==5e-7, 16 samples per prompt, 8 GPUs (H100) and a local mini-batch size of 32. Training took on average 1 day for 2000 steps. Base models with a reasoning chat template are trained with a max token length of 10240, a beta of 0, and a temperature of 1. We report the prompt-level loose accuracy for both IFEval and IFBench. See App. G for more evaluation details.
@@ 4.1 Training on Multiple Constraints
Figure 4: Training on IFTrain (ood) + n constraints (in-domain) from IFEval. (Tülu-DPO policy and Qwen2.5)
@@ 4.3 Changing the range of constraint variables
Most of the constraint templates contain variables which can be instantiated with different values. In your response, all lowercase words should appear at most N times., for example, has the variable $N$ which could in theory be any number. For both the IFEval and the IFBench benchmarks, variables are instantiated for each instruction instance from a fixed range of options. Another type of generalization to assess is whether a model trained on the same constraints, but different variable ranges, can still apply the skill to unseen ranges. To evaluate this, we chose variable ranges for training that are disjoint from the test ranges. For the constraint Your response should contain at most num_sentences sentences., for example, we sample a value between 20 and 40 for train and a value between 1 and 20 for test. Specifically we experiment with three different settings: different range, where every train variable is filled from a range that is disjoint from the test range; wider range, where every train variable is filled from a range that includes and extends the test range; same range, where variables are filled with the same train/test range. Fig. 5 shows performance on IFEval for different steps of IF-RLVR training. Training on a different variable range consistently scores lower than the other two setups. Though training on a wider variable range, performs comparably and often even better than training on the same range. This suggests that training on a diverse set of constraint variables improves generalization for in-domain constraints performance.
@@ 4.4 Removing Categories
Most verifiable constraints fall under broader constraint type categories. IFEval, for example, has 9 different categories, such as length constraints or detectable format. To investigate how training on a set of categories generalizes to unseen categories, we experiment with training on a leave-one-out set of categories, iteratively removing one of the following classes: change cases, detectable format, length constraints, keywords.
Removing the constraints from the length and the keywords categories harms IFEval performance the most, while removing constraints from the change cases and detectable format categories barely affect performance with the model achieving an IFEval accuracy of 89.65 (see Fig. 6).
@@ 4.5 Teaching the basic units of precise IF
We designed the new training constraints so that they would cover IF skills models are currently lacking in, such as copying from the input, counting, and formatting. We find that GRPO training on our new constraints shows targeted improvements in all these areas. As seen in Table 2, our final two models (from base vs. from instruct) improve over the base model (in this case a DPO trained model) in all categories, such as in counting, inserting the right amount of keywords, formatting and in following length constraints. Most of the IFEval categories seem saturated with IF-RLVR training (performance >90>90), while the IFBench categories leave room for improvement, such as the words and sentence categories.
Comparing our final models to other post-trained (SFT, DPO, RLVR) models in Table 3, we find that our approach and model results in the best in- and out-of-domain instruction following performance on both IFEval and IFBench. We also see that targeted RLVR training for IF slightly harms other downstream evaluations, such as AlpacaEval 2, while staying comparable on others, such as GSM8K, MMLU and BBH. We therefore suggest, for future work, to investigate how to combine precise IF RLVR with other types of rewards for other tasks such as math or chat. Note that the performance on other benchmarks is low for IF-RLVR Qwen2.5, as this base model policy has not been post-trained.
@@ 4.6 DPO
We generate prompts using the approach described in Section 3, with up to 5 constraints per prompt, and sample completions from 5 different models (Tülu-3-70B, Qwen-72b, Llama-3.1-405b, Llama3-8b, Yi-34B-Chat). For each constraint in a prompt a completion is verified on whether it fulfills the constraint. We construct preference pairs by sampling, for each instruction, a completion that satisfies all constraints (chosen) and a completion that does not satisfy at least one constraint (rejected). Table 4, shows how most models struggle with precisely following a lot of our prompts, and even when relaxing the requirement and counting cases where models got not more than 1 constraint wrong, the percentages are less than half. The model that stands out is Qwen-72B, where 26% of its completions adhere to all constraints in an instruction and 100% of its completions get only 1 out of up to 5 constraints wrong. Out of all the instances that models get completely correct, 54% have only 1 constraint and only 2% have 5 constraints. We also find that most LLMs get the same easy instances right and the same hard instances wrong, which makes the creation of preference pairs more difficult. This further highlights the benefits of RLVR, where we can get ground truth signal for both easy and hard prompts, with an unlimited amount of constraints. Many of the rejected completions do not contain a single passing constraint out of 5 constraints (46%). We end up with a set of prompts and chosen/rejected pairs, called strict, where chosen completions get all constraints correct (31751 prompts).
@@ 4.7 RLVR for IF from Base
We compare IF-RLVR training using policy models that have gone through post-training already (SFT and DPO), with IF-RLVR training from base models, as it has been shown that it is also possible to perform RLVR training on a base model, leading to good math and reasoning performance shao2024deepseekmath; wang2025reinforcement; yang2025qwen3. Specifically, we RLVR train three base models, llama3.1-8b, Qwen2.5-7B and Qwen3-8B, and their instruct counterparts. To encourage reasoning we use a special chat template (see Appendix F). The models are trained with a max token length of 10240, a beta of 0, and a temp. of 1.
@@ 4.8 RLVR for Multi-turn IF
We experiment with comparing single-turn RLVR runs (IF-RLVR-single) to multi-turn RLVR training (IF-RLVR-multi), and to training on a mix of both types (IF-RLVR-mix). We run experiments on Qwen2.7-7b base and instruct. IF-RLVR-multi mostly leads to an improved performance on the multi-turn setup of IFBench, compared to IF-RLVR-single, without achieving comparable single-turn performance (Table 7). IF-RLVR-mix improves single-turn performance, sometimes beyond improvements from IF-RLVR-single, while reaching a comparable multi-turn performance.
Table 7: Training on single-turn data, multi-turn data, and a mix. Evaluated on IFEval (IFE.) constraints and IFBench (IFB.) constraints.
trained on single-turn trained on multi-turn trained on a mix
single-turn multi-turn single-turn multi-turn single-turn multi-turn
@@ 5 Reward Hacking and the Instruction Hierarchy
To understand the trade-off between challenging constraints and general response quality we contrast IFBench accuracy with a general LLM-as-a-judge prompt-completion rating (zheng2023judging). Using GPT4.1 as the judge, we prompt [^3] See Appendix C for details. it to score how well a completion answers a prompt without the constraint. We score the IFEval and IFBench completions from our RLVR trained model and from the base policy before RLVR training. Completions from the base policy are scored higher by the LLM-as-judge than completions from the IF-RLVR trained model, for both IFEval and IFBench prompts with constraints removed, with the base policy completions receiving an average score of 7 out of 10 and the IF-RLVR trained completions an average score of 6.4 (see Figure 8 in the Appendix for a visualization). This indicates that the base policy models are better at following general instructions, while IF-RLVR trained models are better at following the constraints. The verifiable accuracy, though, is on average higher for the model that went through RLVR training. The verifiable rewards teach the model to prioritize the constraints and in the following section we propose an approach to soften this preference. We also analyze IFBench completions of existing models: claude-3-7-sonnet, Gemini, Qwen2.5-72B-Instruct, and tulu3-70B. We find that tulu3-70B’s IF accuracy correlates the most with the LLM-as-judge scores and that Qwen2.5-72B-Instruct’s IF accuracy is the most negatively correlated with LLM-as-judge scores, out of this set of models. This indicates that Qwen2.5 tends to focus on the constraints rather than the general instruction. We propose adding a reward model signal to the verifiable reward to mitigate reward hacking (App. E).